# Spatial distribution of organics in the Horsehead nebula: signposts of chemistry driven by atomic carbon

### v2

@@ 2.2 IRAM 30m observations
2.3 Interferometric and single-dish data combination
We combined the IRAM 30m and ALMA-7m to produce a clean image containing all spatial information from 0 to $50$ m, i.e., equivalent to a total power telescope of 5050m-diameter. The IRAM 30m map is deconvolved from the $30$ m beam in the Fourier plane before multiplication by the ALMA-7m primary beam in the image plane. After a last Fourier transform, pseudo-visibilities were sampled between 0 and $7$ m. These visibilities were then merged with the ALMA-7m interferometric observations (see Rodriguez-Fernandez et al. 2008, for details). Standard imaging, CLEAN deconvolution, and primary beam correction steps were then applied. The resulting data cube was then scaled from Jy beam-1 to $T_{\mathrm{mb}}$ temperature scale using the synthesized beam size. All these processes happened in the GILDAS/MAPPING software.
The observational parameters of the final combined (hereinafter ALMA-7m+30m) images are summarized in Table 1. The channel rms is calculated using the central 50% of the pixels in the first and last five (line-free) channels of the data cubes. To compare the spatial distribution of the different molecular tracers, we collapse the line emission into velocity-integrated intensity, or zeroth-moment, maps using the CASA task immoments. The zeroth-moment rms is estimated as $\sigma_{\mathrm{rms}}\approx\sigma_{\mathrm{chan}}\sqrt{N_{\mathrm{chan}}}\Delta v$, where $\sigma_{\mathrm{chan}}$ is the channel rms, $N_{\mathrm{chan}}$ is the number of channels used to compute the zeroth-moment, and $\Delta v$ is the channel spacing in velocity.
2.4 Legacy data
The analysis of the ALMA-7m+30m data was complemented with PdBI observations of formaldehyde (H2CO) $J_{K_{a},K_{c}}=2_{02}-1_{01}$ and methanol (CH3OH) $J_{K}=3_{-1}-2_{-1}$ (E symmetry) rotational transitions imaged at $\theta_{\mathrm{beam}}\sim 6^{\prime\prime}$. All the observational details of these legacy data can be found in Guzmán et al. (2013). In order to make the two data sets comparable, the PdBI data was rotated by 14, and the angular resolution was downgraded to $\theta_{\mathrm{beam}}\sim 15^{\prime\prime}$ to match the ALMA-7m+30m data using the CASA tools rotate and convolve2d, respectively. Afterward, the spatial axes were regridded using the CASA task imregrid to match the number of pixels per beam. Throughout this procedure, the CH3CN image was used as a template. Thus, for all subsequent analysis, the angular resolution of the H2CO and CH3OH images is exactly the same as for CH3CN (see Table 1).
@@ 3 Results
This section presents the main findings from our combined ALMA-7m+30m observations. We describe the spatial distribution of simple and complex organics in the Horsehead nebula, based on integrated intensity maps and their corresponding horizontal profiles at key vertical positions. Next, we detail the radiative transfer modeling used to constrain the physical conditions and derive column density profiles for each organic species. Finally, we compare our results with previous studies in the same source and discuss the differences and similarities among the species analyzed.
@@ 3.1 Spatial distribution of molecules
Similar to previous works (e.g., Guzmán et al. 2015; Hernández-Vera et al. 2023), we extracted horizontal cuts from the maps to quantify the distribution of the line emission as a function of $\delta x$, particularly at the $\delta y$ position of the PDR and dense core. In this case, we averaged the integrated line intensity over $10^{\prime\prime}$ in the $\delta y$ direction since it roughly matches the beam of the observations in that direction (see right panels, Fig. 3). Hereinafter, we decided to include the convolved maps of H2CO and CH3OH in the analysis to compare the emission of these smaller species with the larger COMs presented here. Moreover, they can be essential precursors of more complex species (Öberg et al. 2009; López-Sepulcre et al. 2024).
@@ 3.2 Radiative transfer modeling
We have performed radiative transfer calculations to model the line intensities shown in Fig. 3. Typically, the usual way of doing this is through a rotational diagram analysis (Goldsmith and Langer 1999), which assumes optically thin emission and level populations following a Boltzmann distribution characterized by a single excitation temperature ($T_{\mathrm{ex}}$) and the total column density of the molecule ($N$). Under local thermodynamic equilibrium (LTE), $T_{\mathrm{ex}}$ provides a good approximation of the gas kinetic temperature ($T_{\mathrm{kin}}$). However, previous studies have shown that complex molecules in the Horsehead edge are, in general, sub-thermally excited ($T_{\mathrm{ex}}<T_{\mathrm{kin}}$; see Guzmán et al. 2014, and references therein). Therefore, to obtain more accurate results, it is always desirable to take a non-LTE approach, in which both collisional and radiative de-excitation must be considered.
We used the non-LTE radiative transfer code RADEX (van der Tak et al. 2007). More specifically, we ran our analyses using the Python-wrapped version SpectralRadex [^7] https://github.com/uclchem/SpectralRadex (Holdship et al. 2021). Taking this approach, the modeled emission of a specific molecule (in units of K km s-1) will depend on three free parameters: the gas density of the collisional partners, typically H2 ($n_{\mathrm{H_{2}}}$) [^8] We neglect the possible influence of H, He, and e- as collisional partners., the gas temperature ($T_{\mathrm{kin}}$), and the total column density of the molecule ($N$). Other parameters, such as the background temperature ($T_{\mathrm{bg}}$) and the linewidth ($\Delta v$), can also be varied. For simplicity, and because they are not expected to vary significantly across the analyzed region, we kept them fixed at $T_{\mathrm{bg}}=2.73$ K (CMB temperature) and $\Delta v=0.6$ km s-1, as supported by previous studies (Gratier et al. 2013; Guzmán et al. 2014).
@@ 3.2.1 C17O-to-H2 conversion factor and $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})$ ratio
Although the collisional rates of C17O with H2 are available (Yang et al. 2010), we decided to model the emission using the LTE approach. This is a reasonable assumption, given that the critical density of the C17O J=10J=1-0 transition is generally $n_{\mathrm{crit}}<n_{\mathrm{H_{2}}}$ (see Table 3) for the typical gas densities ($n_{\mathrm{H_{2}}}$ $\sim 10^{4}-10^{5}$ cm-3) and temperatures ($T_{\mathrm{kin}}$ $\lesssim 100$ K) observed in the inner layers of the Horsehead nebula (Habart et al. 2005; Pety et al. 2005). Furthermore, we modeled the spectral lines directly, rather than the integrated intensity, to properly account for the hyperfine structure of the emission. Following the procedure described in Hernández-Vera et al. (2024), the line intensity ($I_{\nu}$) in LTE can be expressed in terms of $N$, $T_{\mathrm{ex}}$, and $\Delta v$. However, since we only had one C17O transition observed with ALMA-7m+30m, it was not possible to simultaneously constrain all three parameters.
@@
Hence, with the excitation temperature fixed at $T_{\mathrm{ex}}=15$ K, we repeated the fitting procedure but now for the ALMA-7m+30m data, leaving $N$ and $\Delta v$ as free parameters. This approach allows us to determine NN(C17O)—needed to estimate $n_{\mathrm{H_{2}}}$—while $\Delta v$ is constrained by fitting the hyperfine components of the observed transition. The best-fit models for the PDR and dense core positions are shown in Fig. 8. Defining the C17O-to-H2 conversion factor as $X{{}_{\mathrm{C^{17}O}}}=N(\mathrm{H_{2}})/N(\mathrm{C^{17}O})$, we combined the $N(\mathrm{H_{2}})$ estimates for the PDR and dense core from Guzmán et al. (2014) with our $N(\mathrm{C^{17}O})$ measurements to derive empirical conversion factors for the Horsehead nebula. The resulting values are $X_{\mathrm{C^{17}O}}=(4.0\pm 0.2)\times 10^{6}$ for the PDR and $X_{\mathrm{C^{17}O}}=(5.9\pm 0.3)\times 10^{6}$ for the dense core.
@@ 3.2.2 Gas density based on C17O observations
Once the $X_{\mathrm{C^{17}O}}$ conversion factor was established, we extended the same fitting procedure of C17O ALMA-7m+30m data to all other $\delta x$ positions at the $\delta y$ locations of the PDR and dense core, using the range shown in Fig. 3. This yielded $N(\mathrm{C^{17}O})$ profiles that we converted into $n_{\mathrm{H_{2}}}$ profiles. The advantage of this method is that, by using ALMA-7m+30m data, we obtain empirical gas density profiles at the same angular resolution as the other molecular line observations, thus enabling more accurate column density determinations for the organic species analyzed in this work.
Assuming a typical cloud depth along the line of sight of $\ell_{\mathrm{depth}}\simeq 0.1$ pc (Habart et al. 2005; Lis et al. 2025), and accounting for the fact that the empirical $X_{\mathrm{C^{17}O}}$ value varies between the PDR and dense core, we define the conversion from NN(C17O) to $n_{\mathrm{H_{2}}}$ as follows:
@@
The top panel of Fig. 4 shows the resulting $n_{\mathrm{H_{2}}}$ profiles for the PDR and dense core vertical positions, along with the profile modeled by Habart et al. (2005) convolved with a $15^{\prime\prime}$ FWHM Gaussian to facilitate comparison with the ALMA-7m+30m data. Both profiles confirm the presence of a steep density gradient, with typical values of $n_{\mathrm{H_{2}}}\sim 10^{4}-10^{5}$ cm-3. Although lower densities in the surface layers of the cloud are also plausible ($n_{\mathrm{H_{2}}}\lesssim 10^{3}$ cm-3, Hernández-Vera et al. 2023), non-LTE effects and beam pickup from colder and denser regions could still bias the inferred C17O column densities upward, and thus the $n_{\mathrm{H_{2}}}$ values derived at the edge of the cloud.
@@ 3.2.3 Gas temperature based on CH3CN observations
As described at the beginning of Sect. 3.2, we used SpectralRadex to model the integrated line intensity of the three CH3CN J=65J=6-5 $(K=0,1,\mathrm{and}\>2)$ transitions simultaneously. In this case, we decided to employ the non-LTE approach since Gratier et al. (2013) demonstrated that CH3CN emission at the Horsehead PDR is sub-thermally excited. The collisional rates were taken from Ben Khalifa et al. (2023), and we assumed a ratio of $\mathrm{E/A}=1$ between the two non-interacting torsional substates. Using the $n_{\mathrm{H_{2}}}$ values derived from C17O as input, the modeled CH3CN intensities only depend on $T_{\mathrm{kin}}$ and $N(\mathrm{CH_{3}CN})$ as free parameters. The best-fit integrated intensity models are shown in Fig. 9. The resulting $T_{\mathrm{kin}}$ profiles for the PDR and dense core vertical positions are shown in the middle panel of Fig. 4, whereas the $N(\mathrm{CH_{3}CN})$ profiles are shown in Fig. 5 and Fig. 10.
@@ 3.2.4 Gas thermal pressure
Having derived empirical profiles for the gas density and kinetic temperature, we computed the corresponding thermal pressure, defined as $P_{\mathrm{th}}=n_{\mathrm{H}}T_{\mathrm{kin}}$, where $n_{\mathrm{H}}=n(\mathrm{H})+2n(\mathrm{H_{2}})\sim 2n_{\mathrm{H_{2}}}$ corresponds to the total hydrogen nuclei density. The resulting $P_{\mathrm{th}}$ profiles for the vertical positions of the PDR and dense core are shown in the bottom panel of Fig. 4, together with the range of values determined from high angular resolution observations by Hernández-Vera et al. (2023) [^9] Note that in Hernández-Vera et al. (2023) the factor of 2 in the definition of $n_{\mathrm{H}}\sim 2n_{\mathrm{H_{2}}}$ is missing. As a result, and for consistency with previous work, the range of values used for comparison here is $P_{\mathrm{th}}=(4.6$$8.0)\times 10^{6}$ K cm-3..
@@
Figure 5: Column density profiles derived at the $\delta y$ position of the PDR from the radiative transfer modeling of the O-bearing (left panel) and N-bearing (right panel) molecules analyzed in this work. For each color, the solid line depicts the best-fit values, whose uncertainties are represented by the shaded areas of the same color. The red vertical line, the dotted vertical lines, and the horizontal gray bar are the same as in the left panels of Fig. 3. The same profiles but at the $\delta y$ position of the dense core are shown in Fig. 10.
@@ 3.2.5 Column density profiles
The integrated line emission of the remaining molecules was modeled to obtain column density profiles and further investigate the formation pathways of the different chemical species. Although it was mentioned that complex molecules in the Horsehead nebula are sub-thermally excited (i.e., $n_{\mathrm{crit}}<n_{\mathrm{H_{2}}}$, see Table 3), we cannot use the RADEX non-LTE approach for all the species presented in this work since not all have collisional rates available in the literature.
Thus, for those species with available rates, such as H2CO (Wiesenfeld and Faure 2013), CH3OH (Rabli and Flower 2010), HNCO (Sahnoun et al. 2018), and HC3N (Faure et al. 2016), we performed a similar procedure as CH3CN but using the $n_{\mathrm{H_{2}}}$ and $T_{\mathrm{kin}}$ profiles derived in the previous sections as input. Instead, for CH2CO and CH3CHO, we used the rotational diagram approach by assuming a Boltzmann distribution of the rotational levels characterized by a single $T_{\mathrm{ex}}$ lower than $T_{\mathrm{kin}}$, consistent with a sub-thermal excitation regime. Under this assumption, the column density in the upper state of a given transition ($N_{u}$) is given by:
@@
where $N$ is the total column density, $Q$ is the partition function of the molecule taken from the JPL database (Pickett et al. 1998), and $E_{u}$ with $g_{u}$ are the energy and the degeneracy associated to the upper-state level of the corresponding transition, respectively (see Table 3). On the other hand, assuming optically thin emission, $N_{u}$ is related to the integrated line intensity ($I=\int I_{\nu}d\nu$, in K km s-1) through the equation:
@@
Figure 5 shows the derived column density profiles for the PDR vertical position, while the corresponding profiles for the dense core vertical position are presented in Fig. 10. Since the profiles at both positions are very similar, we will refer only to Fig. 5 in the subsequent analysis, with the understanding that the same conclusions apply to the dense core vertical position. The associated uncertainties correspond to the statistical propagation of observational errors, without including systematic uncertainties on the fixed parameters, as this would require exploring a multidimensional parameter space in density and temperature, which is beyond the scope of this work. We checked the consistency of our results by comparing them with previous efforts. Since this is the first time that the column density of these molecules is spatially imaged, the only benchmarks that we have are the values derived at the PDR and dense core positions from prior single-dish observations, except for HNCO for which there are no results reported in the literature.
@@
For the molecules observed with the PdBI, our results are consistent within the uncertainties with the values determined by Guzmán et al. (2013) using the IRAM 30m spectra only, confirming that the emission of H2CO and CH3OH is only marginally diluted. Instead, for the combined ALMA-7m+30m maps, most of our column densities are higher than previous results from the WHISPER line survey (Gratier et al. 2013; Guzmán et al. 2014), specifically by a factor of $\sim 3-4$ in the dense core and a factor of $\sim 3-6$ in the PDR. Beam dilution effects can explain the above since, from the maps shown in Fig. 2, we can see that we begin to resolve filamentary and clumpy structures with angular widths on the order of $\theta_{\mathrm{source}}\sim 10-15^{\prime\prime}$ that probably were diluted by the large beam size reached by WHISPER ($\theta_{\mathrm{beam}}\gtrsim 27^{\prime\prime}$, Guzmán et al. 2014).
@@
For CH3CN, at the PDR position we found a column density $\sim 0.5$ times lower than that reported by Gratier et al. (2013) using WHISPER data. However, Gratier et al. (2013) already accounted for beam dilution in their radiative transfer modeling, which explains why we find a similar column density with the new higher-angular resolution observations. At the dense core, in contrast, we find a higher CH3CN column density (by a factor of $\sim 3$) compared to Gratier et al. (2013). This discrepancy is likely due to the choice of collisional rates: Gratier et al. (2013) adopted the rates from Green (1986), whereas Ben Khalifa et al. (2023) recently showed that those rates can be severely underestimated for some transitions, especially at low temperatures ($T_{\mathrm{kin}}=20$ K) such as those found in the dense core. Indeed, when we repeated our analysis using the older rates from Green (1986), our derived $N$ values matched those reported by Gratier et al. (2013). In the PDR, on the other hand, the choice of collisional coefficients has a less significant effect.

### v3

@@ 3 Results
This section presents the main findings from our combined ALMA 7m+30m observations. We describe the spatial distribution of simple and complex organics in the Horsehead nebula, based on integrated intensity maps and their corresponding horizontal profiles at key vertical positions. Next, we detail the radiative transfer modeling used to constrain the physical conditions and derive column density profiles for each organic species. Finally, we compare our results with previous studies in the same source and discuss the differences and similarities among the species analyzed.
@@ 3.1 Spatial distribution of molecules
Similar to previous works (e.g., Guzmán et al. 2015; Hernández-Vera et al. 2023), we extracted horizontal cuts from the maps to quantify the distribution of the line emission as a function of $\delta x$, particularly at the $\delta y$ position of the PDR and dense core. In this case, we averaged the integrated line intensity over $10^{\prime\prime}$ in the $\delta y$ direction since it roughly matches the beam of the observations in that direction (see right panels, Fig. 3). Hereafter, we decided to include the convolved maps of H2CO and CH3OH in the analysis to compare the emission of these smaller species with the larger COMs presented here. Moreover, they can be essential precursors of more complex species (Öberg et al. 2009; López-Sepulcre et al. 2024).
@@ 3.2 Radiative transfer modeling
We have performed radiative transfer calculations to model the line intensities shown in Fig. 3. Typically, the usual way of doing this is through a rotational diagram analysis (Goldsmith and Langer 1999), which assumes optically thin emission and level populations following a Boltzmann distribution characterized by a single excitation temperature ($T_{\mathrm{ex}}$) and the total column density of the molecule ($N$). Under local thermodynamic equilibrium (LTE), $T_{\mathrm{ex}}$ provides a good approximation of the gas kinetic temperature ($T_{\mathrm{kin}}$). However, previous studies have shown that complex molecules in the Horsehead edge are, in general, subthermally excited ($T_{\mathrm{ex}}<T_{\mathrm{kin}}$; see Guzmán et al. 2014, and references therein). Therefore, to obtain more accurate results, it is always desirable to take a non-LTE approach, in which both collisional and radiative de-excitation must be considered.
We used the non-LTE radiative transfer code RADEX (van der Tak et al. 2007). More specifically, we ran our analyses using the Python-wrapped version SpectralRadex [^7] https://github.com/uclchem/SpectralRadex (Holdship et al. 2021). Taking this approach, the modeled emission of a specific molecule (in units of K km s-1) will depend on three free parameters: the gas density of the collisional partners, typically H2 ($n_{\mathrm{H_{2}}}$), [^8] We neglect the possible influence of H, He, and e- as collisional partners. the gas temperature ($T_{\mathrm{kin}}$), and the total column density of the molecule ($N$). Other parameters, such as the background temperature ($T_{\mathrm{bg}}$) and the linewidth ($\Delta v$), can also be varied. For simplicity, and because they are not expected to vary significantly across the analyzed region, we kept them fixed at $T_{\mathrm{bg}}=2.73$ K (CMB temperature) and $\Delta v=0.6$ km s-1, as supported by previous studies (Gratier et al. 2013; Guzmán et al. 2014).
@@ 3.2.1 C17O-to-H2 conversion factor and $N(\mathrm{C^{18}O})/N(\mathrm{C^{17}O})$ ratio
Although the collisional rates of C17O with H2 are available (Yang et al. 2010), we decided to model the emission using the LTE approach. This is a reasonable assumption, given that the critical density of the C17O J=10J=1-0 transition is generally $n_{\mathrm{crit}}<n_{\mathrm{H_{2}}}$ (see Table 3) for the typical gas densities ($n_{\mathrm{H_{2}}}$ $\sim 10^{4}-10^{5}$ cm-3) and temperatures ($T_{\mathrm{kin}}$ $\lesssim 100$ K) observed in the inner layers of the Horsehead nebula (Habart et al. 2005; Pety et al. 2005). Furthermore, we modeled the spectral lines directly, rather than the integrated intensity, to properly account for the hyperfine structure of the emission. Following the procedure described in Hernández-Vera et al. (2024), the line intensity ($I_{\nu}$) in LTE can be expressed in terms of $N$, $T_{\mathrm{ex}}$, and $\Delta v$. However, since we only had one C17O transition observed with ALMA 7m+30m, it was not possible to simultaneously constrain all three parameters.
@@
Hence, with the excitation temperature fixed at $T_{\mathrm{ex}}=15$ K, we repeated the fitting procedure but now for the ALMA 7m+30m data, leaving $N$ and $\Delta v$ as free parameters. This approach allows us to determine NN(C17O)—needed to estimate $n_{\mathrm{H_{2}}}$—while $\Delta v$ is constrained by fitting the hyperfine components of the observed transition. The best-fit models for the PDR and dense core positions are shown in Fig. 8. Defining the C17O-to-H2 conversion factor as $X{{}_{\mathrm{C^{17}O}}}=N(\mathrm{H_{2}})/N(\mathrm{C^{17}O})$, we combined the $N(\mathrm{H_{2}})$ estimates for the PDR and dense core from Guzmán et al. (2014) with our $N(\mathrm{C^{17}O})$ measurements to derive empirical conversion factors for the Horsehead nebula. The resulting values are $X_{\mathrm{C^{17}O}}=(4.0\pm 0.2)\times 10^{6}$ for the PDR and $X_{\mathrm{C^{17}O}}=(5.9\pm 0.3)\times 10^{6}$ for the dense core.
@@ 3.2.2 Gas density based on C17O observations
Once the $X_{\mathrm{C^{17}O}}$ conversion factor was established, we extended the same fitting procedure of C17O ALMA 7m+30m data to all other $\delta x$ positions at the $\delta y$ locations of the PDR and dense core, using the range shown in Fig. 3. This yielded $N(\mathrm{C^{17}O})$ profiles that we converted into $n_{\mathrm{H_{2}}}$ profiles. The advantage of this method is that, by using ALMA 7m+30m data, we obtain empirical gas density profiles at the same angular resolution as the other molecular line observations, thus enabling more accurate column density determinations for the organic species analyzed in this work.
Assuming a typical cloud depth along the line of sight of $\ell_{\mathrm{depth}}\simeq 0.1$ pc (Habart et al. 2005; Lis et al. 2026), and accounting for the fact that the empirical $X_{\mathrm{C^{17}O}}$ value varies between the PDR and dense core, we defined the conversion from NN(C17O) to $n_{\mathrm{H_{2}}}$ as
@@
The top panel of Fig. 4 shows the resulting $n_{\mathrm{H_{2}}}$ profiles for the PDR and dense core vertical positions, along with the profile modeled by Habart et al. (2005) convolved with a $15^{\prime\prime}$ full width at half maximum (FWHM) Gaussian to facilitate comparison with the ALMA 7m+30m data. Both profiles confirm the presence of a steep density gradient, with typical values of $n_{\mathrm{H_{2}}}\sim 10^{4}-10^{5}$ cm-3. Although lower densities in the surface layers of the cloud are also plausible ($n_{\mathrm{H_{2}}}\lesssim 10^{3}$ cm-3, Hernández-Vera et al. 2023), non-LTE effects and beam pickup from colder and denser regions could still bias the inferred C17O column densities upward, and thus the $n_{\mathrm{H_{2}}}$ values derived at the edge of the cloud.
@@ 3.2.3 Gas temperature based on CH3CN observations
As described at the beginning of Sect. 3.2, we used SpectralRadex to model the integrated line intensity of the three CH3CN J=65J=6-5 $(K=0,1,\mathrm{and}\>2)$ transitions simultaneously. In this case, we decided to employ the non-LTE approach since Gratier et al. (2013) demonstrated that CH3CN emission at the Horsehead PDR is subthermally excited. The collisional rates were taken from Ben Khalifa et al. (2023), and we assumed a ratio of $\mathrm{E/A}=1$ between the two noninteracting torsional substates. Using the $n_{\mathrm{H_{2}}}$ values derived from C17O as input, the modeled CH3CN intensities only depend on $T_{\mathrm{kin}}$ and $N(\mathrm{CH_{3}CN})$ as free parameters. The best-fit integrated intensity models are shown in Fig. 9. The resulting $T_{\mathrm{kin}}$ profiles for the PDR and dense core vertical positions are shown in the middle panel of Fig. 4, whereas the $N(\mathrm{CH_{3}CN})$ profiles are shown in Fig. 5 and Fig. 10.
@@ 3.2.4 Gas thermal pressure
Having derived empirical profiles for the gas density and kinetic temperature, we computed the corresponding thermal pressure, defined as $P_{\mathrm{th}}=n_{\mathrm{H}}T_{\mathrm{kin}}$, where $n_{\mathrm{H}}=n(\mathrm{H})+2n(\mathrm{H_{2}})\sim 2n_{\mathrm{H_{2}}}$ corresponds to the total hydrogen nuclei density. The resulting $P_{\mathrm{th}}$ profiles for the vertical positions of the PDR and dense core are shown in the bottom panel of Fig. 4, together with the range of values determined from high angular resolution observations by Hernández-Vera et al. (2023). [^9] In Hernández-Vera et al. (2023) the factor of 2 in the definition of $n_{\mathrm{H}}\sim 2n_{\mathrm{H_{2}}}$ is missing. As a result, and for consistency with previous work, the range of values used for comparison here is $P_{\mathrm{th}}=(4.6$$8.0)\times 10^{6}$ K cm-3.
@@
Figure 5: Column density profiles derived at the $\delta y$ position of the PDR from the radiative transfer modeling of the O-bearing (left panel) and N-bearing (right panel) molecules analyzed in this work. For each color the solid line depicts the best-fit values, whose uncertainties are represented by the shaded areas of the same color. The red vertical line, the dotted vertical lines, and the horizontal gray bar are the same as in the left panels of Fig. 3. The same profiles, but at the $\delta y$ position of the dense core, are shown in Fig. 10.
@@ 3.2.5 Column density profiles
The integrated line emission of the remaining molecules was modeled to obtain column density profiles and further investigate the formation pathways of the different chemical species. Although it was mentioned that complex molecules in the Horsehead nebula are subthermally excited (i.e., $n_{\mathrm{crit}}<n_{\mathrm{H_{2}}}$, see Table 3), we cannot use the RADEX non-LTE approach for all the species presented in this work since not all have collisional rates available in the literature.
Thus, for those species with available rates, such as H2CO (Wiesenfeld and Faure 2013), CH3OH (Rabli and Flower 2010), HNCO (Sahnoun et al. 2018), and HC3N (Faure et al. 2016), we performed a similar procedure as CH3CN but using the $n_{\mathrm{H_{2}}}$ and $T_{\mathrm{kin}}$ profiles derived in the previous sections as input. Instead, for CH2CO and CH3CHO, we used the rotational diagram approach by assuming a Boltzmann distribution of the rotational levels characterized by a single $T_{\mathrm{ex}}$ lower than $T_{\mathrm{kin}}$, consistent with a subthermal excitation regime. Under this assumption, the column density in the upper state of a given transition ($N_{u}$) is given by
@@
where $N$ is the total column density, $Q$ is the partition function of the molecule taken from the JPL database (Pickett et al. 1998), and $E_{u}$ with $g_{u}$ are the energy and the degeneracy associated with the upper-state level of the corresponding transition, respectively (see Table 3). On the other hand, assuming optically thin emission, $N_{u}$ is related to the integrated line intensity ($I=\int I_{\nu}d\nu$, in K km s-1) through the equation
@@
Figure 5 shows the derived column density profiles for the PDR vertical position, while the corresponding profiles for the dense core vertical position are presented in Fig. 10. Since the profiles at both positions are very similar, we refer only to Fig. 5 in the subsequent analysis, with the understanding that the same conclusions apply to the dense core vertical position. The associated uncertainties correspond to the statistical propagation of observational errors, without including systematic uncertainties on the fixed parameters, as this would require exploring a multidimensional parameter space in density and temperature, which is beyond the scope of this work. We checked the consistency of our results by comparing them with previous efforts. Since this is the first time that the column density of these molecules is spatially imaged, the only benchmarks that we have are the values derived at the PDR and dense core positions from prior single-dish observations, except for HNCO for which there are no results reported in the literature.
@@
For the molecules observed with the PdBI, our results are consistent within the uncertainties with the values determined by Guzmán et al. (2013) using the IRAM 30m spectra only, confirming that the emission of H2CO and CH3OH is only marginally diluted. Instead, for the combined ALMA 7m+30m maps, most of our column densities are higher than previous results from the WHISPER line survey (Gratier et al. 2013; Guzmán et al. 2014), specifically by a factor of $\sim 3-4$ in the dense core and a factor of $\sim 3-6$ in the PDR. Beam dilution effects can explain the above since, from the maps shown in Fig. 2, we can see that we begin to resolve filamentary and clumpy structures with angular widths on the order of $\theta_{\mathrm{source}}\sim 10-15^{\prime\prime}$ that probably were diluted by the large beam size reached by WHISPER ($\theta_{\mathrm{beam}}\gtrsim 27^{\prime\prime}$, Guzmán et al. 2014).
@@
For CH3CN, at the PDR position we found a column density $\sim 0.5$ times lower than that reported by Gratier et al. (2013) using WHISPER data. However, Gratier et al. (2013) already accounted for beam dilution in their radiative transfer modeling, which explains why we find a similar column density with the new higher angular resolution observations. At the dense core, in contrast, we find a higher CH3CN column density (by a factor of $\sim 3$) compared to Gratier et al. (2013). This discrepancy is likely due to the choice of collisional rates: Gratier et al. (2013) adopted the rates from Green (1986), whereas Ben Khalifa et al. (2023) recently showed that those rates can be severely underestimated for some transitions, especially at low temperatures ($T_{\mathrm{kin}}=20$ K) such as those found in the dense core. When we repeated our analysis using the older rates from Green (1986), our derived $N$ values matched those reported by Gratier et al. (2013). In the PDR, on the other hand, the choice of collisional coefficients has a less significant effect.